Increasing the average speed of the Pickerington Express Bus by 13 km/h resulted in the 260 km trip taking a hour less than before. What was the original average speed of the bus?
Answer by mananth(16949)

let the original speed of the bus be x
speed increased by 13 km/h , x+13
distance = 260
original time = new time + 1
original time - new time = 1
t=d/r
260/x - 260/(x+13)=1
LCD = x(x+13)
multiply by the LCD
260(x+13) -260x= x(x+13)
260x+3380-260x=x^2+13x
x^2+13x-3380=0
x^2+65x-52x-3380=0
x(x+65)-52(x+65)=0
(x-52)(x+65)=0
x= 52 km/h the original speed
we ignore the negative value x=-65